please help me solve this, its a triangle with a hypotenuse of 12, leg a is x and leg b is (x+5)
Answer by scott8148(6628)

x^2 + (x + 5)^2 = 12^2

2x^2 + 10x + 25 = 144 ___ 2x^2 + 10x - 119 = 0

use the quadratic formula to find x



***************** ANSWER BELOW *****************


x = 5.61 (approx)